### v1

@@ Abstract
We consider a model for explainable AI in which an explanation for a prediction $h(x)=y$ consists of a subset $S^{\prime}$ of the training data (if it exists) such that all classifiers $h^{\prime}\in\mathcal{H}$ that make at most $b$ mistakes on $S^{\prime}$ predict $h^{\prime}(x)=y$. Such a set $S^{\prime}$ serves as a proof that $x$ indeed has label $y$ under the assumption that (1) the true target function $h^{\star}$ belongs to $\mathcal{H}$, and (2) the set $S$ contains at most $b$ noisy or corrupted points. For example, if $b=0$ and $\mathcal{H}$ is the family of linear classifiers in $\mathbb{R}^{d}$, and if $x$ lies inside the convex hull of the positive data points in $S$ (and therefore every consistent linear classifier labels $x$ as positive), then Carathéodory’s theorem states that $x$ in fact lies inside the convex hull of $d+1$ of those points. So, a set $S^{\prime}$ of size $d+1$ could be released as an explanation for a positive prediction, and would serve as a short proof of correctness of the prediction under the assumption of perfect realizability.
In this work, we consider this problem more generally, for general hypothesis classes $\mathcal{H}$ and general values $b\geq 0$. We define the notion of the robust hollow star number of $\mathcal{H}$ (which generalizes the standard hollow star number), and show that it precisely characterizes the worst-case size of the smallest certificate achievable, and analyze its size for natural classes. We also consider worst-case distributional bounds on certificate size, as well as distribution-dependent bounds that we show tightly control the sample size needed to get a certificate for any given test example. In particular, we define a notion of the certificate coefficient $\varepsilon_{x}$ of an example $x$ with respect to a data distribution $\mathcal{D}$ and target function $h^{\star}$, and prove matching upper and lower bounds on sample size as a function of $\varepsilon_{x}$, $b$, and the VC dimension $d$ of $\mathcal{H}$.
@@ 1 Introduction
There has been substantial recent interest in explainable AI, [1, 14, 15, 27, 23]. For example, in a medical setting, if a classifier $h\in\mathcal{H}$ trained on some large dataset $S$ predicts that patient $x$ should get treatment $y$, the patient’s doctor may want an explanation of why. Much of the work in explainable machine learning has focused on decision-tree models, or identifying the most salient features for the prediction made [13, 10, 32]. In this work, we consider an alternative approach: outputting a subset $S^{\prime}$ of the training set $S$ such that all classifiers $h^{\prime}\in\mathcal{H}$ that agree with $S^{\prime}$ (or that make at most $b$ mistakes on $S^{\prime}$) predict $h^{\prime}(x)=y$, if such an $S^{\prime}$ exists. Such a set $S^{\prime}$ would serve as a proof that $x$ indeed has label $y$ under the assumption that (1) the true target function $h^{\star}$ belongs to $\mathcal{H}$, and (2) the set $S$ contains at most $b$ noisy or corrupted points. For example, if $b=0$ and $\mathcal{H}$ is the family of linear classifiers in $\mathbb{R}^{d}$, and if $x$ lies inside the convex hull of the positive data points in $S$ (and hence every consistent linear classifier labels $x$ as positive), then Carathéodory’s theorem states that $x$ in fact lies inside the convex hull of $d+1$ of those points. So, a set $S^{\prime}$ of size $d+1$ could be released as an explanation for a positive prediction, and a proof of its correctness under the assumption of realizability.
We aim to consider such explanations for general families $\mathcal{H}$ and general values of $b$. Our work is inspired by [2] who propose the notion of robustly reliable classifiers that, given an example $x$, output both a label $y$ and a value $b$ with the guarantee that any $h^{\prime}\in\mathcal{H}$ with $h^{\prime}(x)\neq y$ makes strictly more than $b$ mistakes on $S$ (where $b<0$ if $x$ is not in the agreement region of the version space). Our work can be viewed as investigating the shortest proof that can be provided for such a guarantee.
@@ 1.1 Main Contributions
    We formalize the notion of a robust certificate: a subset $S^{\prime}$ of the training data that serves as a proof that a given example $x$ must have label $y$ if the target function belongs to a given class $\mathcal{H}$ and the training set has at most $b$ noisy or corrupted points. To analyze this, we define the notion of the robust hollow star number of $\mathcal{H}$, which generalizes the standard hollow star number [5], and show that it precisely characterizes the worst-case size of the smallest certificate achievable for a class $\mathcal{H}$, and analyze its size for natural classes.
@@
    We examine worst-case distributional bounds on certificate size, showing that in this case, one can achieve tight bounds on certificates achievable from a finite sample in terms of the (standard) hollow star number of [5].
@@ 1.2 Context and Related Work
Explainable ML research largely focuses on decision trees or key predictive features. In these approaches, a certificate for an instance $x$ corresponds to the root-to-leaf path of $x$ in the tree. A widely studied method for explaining a black-box model $f$ involves first learning a decision tree $T$ that closely approximates $f$. Once this surrogate decision tree $T$ is obtained, a certificate for any instance $x$ can be generated by retrieving its associated path in $T$ [13, 10, 32]. [8] proposed algorithms for implicitly learning the surrogate decision trees that approximate the target function, with provable performance guarantees under the uniform distribution. [28] were the first to introduce certificates (which they term anchors) providing high precision by identifying a minimal set of rules that anchor a prediction, ensuring that the output remains stable under small perturbations. They provide an efficient heuristic, based on greedy search, for finding such high-precision certificates. [7] investigates the minimum number of queries required to certify a monotone function’s prediction at a given point. They define as certificate a subset of input coordinates such that fixing these guarantees the function’s value remains unchanged; [20] further investigate this. There is also an increasing interest in hybrid models that are partially interpretable as studied by [16, 17].
Another related line of work considers guarantees in the face of instance-targeted data-poisoning attacks, first considered by [9]. Subsequent work by [30] and [29] demonstrated empirically that such attacks can be highly effective, even when the adversary only adds correctly-labeled data to the training set. These targeted attacks have attracted attention in recent years due to their potential to compromise the trustworthiness of systems [18, 24, 12]. A key concern here is: when can predictions be trusted under such attacks? Most theoretical work on this question has focused on certifying stability of predictions under small changes to the training set [19, 22]. However, recently, [2, 6, 11] consider certifying the actual correctness of predictions, under assumptions of realizability and bounded adversarial power. Our work can be viewed as studying the shortest proof that can be provided for such a guarantee.
@@ 2.1 Notation
The input space is denoted $\mathcal{X}$, and the label space $\{+1,-1\}$. A hypothesis class $\mathcal{H}$ is a subset of $\{-1,+1\}^{\mathcal{X}}$. A sequence $S=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ is realizable by $\mathcal{H}$ if $\exists h\in\mathcal{H}$ such that $\sum_{i=1}^{n}\mathds{1}[h(x_{i})\neq y_{i}]=0$. We use $[n]$ to denote the first $n$ natural numbers, $[n]=\{1,2,\dots,n\}$.
@@ 2.2 Formal Setting
When $b=0$, we refer to the 00-robust agreement region simply as the agreement region.
@@
Our setting assumes that we are given a $b$-robustly realizable training dataset $S$, together with a test point $x$ which satisfies that $(x,y)$ belongs to the $b$-robust agreement region of $S$, for some $y\in\{+1,-1\}$. Namely, $S$ is itself a $b$-robust certificate for $(x,y)$. [^1] The problem of obtaining small-size subsets of the training data that can serve as good-faith certificates for test-time predictions only makes sense if the test point is in the agreement region of the training data to start with. Our primary objective is to analyze and obtain the smallest $S^{\prime}\subseteq S$ such that $S^{\prime}$ continues to be a $b$-robust certificate for $(x,y)$.
@@ 3 Worst Case Bounds on Certificate Size
Consider the class of $d$-dimensional halfspaces passing through the origin, i.e., $\mathcal{X}=\mathbb{R}^{d}$, =$\mathcal{H}=\{x\mapsto\operatorname{sign}(w^{T}x):w\in\mathbb{R}^{d}\}$. Suppose we are given a training dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{H}$. Let $S_{+}=\{x_{i}:i\in[n],y_{i}=+1\}$, $S_{-}=\{-x_{i}:i\in[n],y_{i}=-1\}$, and let $S_{+}\cup S_{-}=\{z_{1},\ldots,z_{n}\}$ for convenience. Note that we negate a negatively labeled $x_{i}$ in $S$ to form $S_{-}$. A test point $(x,+1)$ belongs to the agreement region of $S$ if and only if $x\in\operatorname{Cone}(S_{+}\cup S_{-})$, where $\operatorname{Cone}(\cdot)$ denotes the conic hull. To see this, suppose first that $x\in\operatorname{Cone}(S_{+}\cup S_{-})$. This means that $x=\sum_{i=1}^{n}\alpha_{i}z_{i}$, where $\alpha_{i}\geq 0,\forall i$. Let $w\in\mathbb{R}^{d}$ represent any halfspace that labels all the examples in $S$ correctly. Then, observe that for every $z_{i}\in S_{+}\cup S_{-}$, $w^{T}z_{i}\geq 0$. But this also means that $w^{T}x=\sum_{i=1}^{n}\alpha_{i}\cdot w^{T}z_{i}\geq 0$. Thus, $(x,+1)$ belongs to the agreement region of $S$. In the other direction, suppose that $x\notin\operatorname{Cone}(S_{+}\cup S_{-})$. Then, because $\operatorname{Cone}(S_{+}\cup S_{-})$ is a closed convex set, the separating hyperplane theorem (e.g., see Theorem 1 in [26]) implies the existence of $w\in\mathbb{R}^{d}$, such that $w^{T}x<0$, and $w^{T}y\geq 0$ for $y\in\operatorname{Cone}(S_{+}\cup S_{-})$. In particular, this means that $w$ labels all examples in $S$ correctly. However, $w^{T}x<0$, so $(x,+1)$ is not in the agreement region of $S$.
So, consider a test point $(x,+1)$ that belongs to the agreement region of $S$. By the preceding argument, $x\in\operatorname{Cone}(S_{+}\cup S_{-})$. But then, by Carathéodory’s Theorem, $x$ can be written as a conic combination of at most $d$ points [^2] There are hardness results on computing the “Carathéodory number” of a point set [3]. from $S_{+}\cup S_{-}$, which implies that $(x,+1)$ is in the agreement region of these points. Thus, we can use this subset of $S_{+}\cup S_{-}$ (together with labels), which has size at most $d$, as a certificate for $(x,+1)$.
@@
Consider the class of singletons on a domain of size $n$, i.e., $\mathcal{X}=[n]$, =$\mathcal{H}=\{x\mapsto(-1)^{1-\mathds{1}[x=i]}:i\in[n]\}$. The VC dimension of this class is 1. Suppose that we are given a training dataset $S=\{(1,-1),(2,-1),\ldots,(n-1,-1)\}$. Observe that the test point $(n,+1)$ is in the agreement region of $S$, because the only hypothesis in $\mathcal{H}$ that labels $S$ in the given way labels $n$ positively. Note however, that the test point is not in the agreement region of any proper subset of $S$. We must therefore necessarily provide all the $n-1$ points in $S$ to certify a positive label on $n$.
The instance in the above example has a curious property: as long as there remains a single point in the domain whose label we haven’t observed, it is not possible to completely determine the label of the test point. This, however, is the defining property of another combinatorial quantity in learning theory, known as the hollow star number [5]. For example, the hollow star number is known to lower bound the sample complexity of proper PAC learners [5, Theorem 10]. Indeed, as we show ahead, the hollow star number ends up also being the relevant quantity that characterizes worst-case minimum certificate sizes. First, we define a slightly more generalized version of the hollow star number, which allows us to handle $b\geq 0$ corruptions.
@@
A sequence $T=\{(x_{1},y_{1}),\dots,(x_{s{b}},y_{s{b}})\}$ is a $b$-robust hollow star for $\mathcal{H}$, if there exist integer weights $w_{1},\dots,w_{s{b}}$ and $i^{*}\in[s{b}]$ such that:
@@
    For any $i\in[s{b}]$, $T\setminus\{(x_{i},y_{i},w_{i})\}$ is $b$-robustly realizable by $\mathcal{H}$.
The size $s{b}$ of the largest $b$-robust hollow star is the $b$-robust hollow star number of $\mathcal{H}$.
@@
We observe that setting $b=0$ in the above definition recovers the standard definition (e.g., Definition 3 in [5]) of the hollow star number. We also note that repeats are allowed in the sequence $T$ (i.e., there might be $i\neq j$ where $x_{i}=x_{j}$).
@@
For $b\geq 0$, let $s{b}$ be the $b$-robust hollow star number of $\mathcal{H}$. Then, $s{b}\geq(b+1)(s_{0}-1)+1$.
@@
Let $\mathcal{H}$ be a hypothesis class that has $b$-robust hollow star number $s{b}$, $S$ be a training dataset that is $b$-robustly realizable by $\mathcal{H}$, and $x$ be a test point such that for some $y\in\{-1,1\}$, $(x,y)$ is in the $b$-robust agreement region of $S$. Then, there exists a $b$-robust certificate $S^{\prime}\subseteq S$ for $(x,y)$ that has size at most $s{b}-1$. Furthermore, there exists a training dataset $S$ of size $s{b}-1$ that is $b$-robustly realizable by $\mathcal{H}$, test point $x$ and test label $y\in\{-1,1\}$ which satisfy that $(x,y)$ belongs to the $b$-robust agreement region of $S$, such that no proper subsequence of $S$ is a $b$-robust certificate for $(x,y)$.
@@
Consider $S^{\prime}\subseteq S$ to be the smallest subset of $S$ that is a $b$-robust certificate for $(x,y)$—let $S^{\prime}=\{(x_{1},y_{1}),\dots,(x_{k},y_{k})\}$. Note that $S^{\prime}$ is a minimal $b$-robust certificate for $(x,y)$. We argue that $\{(x_{1},y_{1}),\dots,(x_{k},y_{k}),(x,1-y)\}$ is a $b$-robust hollow star, which implies $k+1\leq s_{b}$.
Consider the weighted sequence $T=\{(x_{1},y_{1},1),\dots,(x_{k},y_{k},1),(x,1-y,b+1)\}$. Now consider any $h\in\mathcal{H}$. If $h(x)=y$, then the weighted error of $h$ on $x$ is $b+1$, and $h$ does not $b$-robustly realize $T$. Otherwise, we have that $h(x)=1-y$. In this case, it must be the case that $\sum_{i=1}^{k}\mathds{1}[h(x_{i})\neq y_{i}]>b$, otherwise $S^{\prime}$ would not be a $b$-robust certificate for $(x,y)$. Summarily, the weighted sequence $T$ is not $b$-robustly realizable by any $h\in\mathcal{H}$.
Now, consider $T^{\prime}=T\setminus\{(x,1-y,b+1)\}$. Since $S$ is $b$-robustly realizable by $\mathcal{H}$, so is $S^{\prime}$, which implies that the weighted sequence $T^{\prime}$ is also $b$-robustly realizable by $\mathcal{H}$. Now, consider $T^{\prime}=T\setminus\{(x_{i},y_{i},1)\}$ for any $i\in[k]$. Then, since $S^{\prime}$ is a minimal $b$-robust certificate for $(x,y)$, there must exist some $h\in\mathcal{H}$ such that $\sum_{j=1,j\neq i}^{k}\mathds{1}[h(x_{j})\neq y_{j}]\leq b$ and $h(x)=1-y$. Otherwise, $S^{\prime}\setminus\{(x_{i},y_{i})\}$ would be a smaller $b$-robust certificate for $(x,y)$. Consequently, such an $h$ $b$-robustly realizes the weighted sequence $T^{\prime}$, and we are done.

### v4

@@ Abstract
We consider a model for explainable AI in which an explanation for a prediction $h(x)=y$ consists of a subset $S^{\prime}$ of the training data (if it exists) such that all classifiers $h^{\prime}\in\mathcal{H}$ that make at most $b$ mistakes on $S^{\prime}$ predict $h^{\prime}(x)=y$. Such a set $S^{\prime}$ serves as a proof that $x$ indeed has label $y$ under the assumption that (1) the true target function $h^{\star}$ belongs to $\mathcal{H}$, and (2) the set $S$ contains at most $b$ noisy or corrupted points. For example, if $b=0$ and $\mathcal{H}$ is the family of linear classifiers in $\mathbb{R}^{d}$, and if $x$ lies inside the convex hull of the positive data points in $S$ (and therefore every consistent linear classifier labels $x$ as positive), then Carathéodory’s theorem states that $x$ in fact lies inside the convex hull of $d+1$ of those points. So, a set $S^{\prime}$ of size $d+1$ could be released as an explanation for a positive prediction, and would serve as a short proof of correctness of the prediction under the assumption of perfect realizability. In this work, we consider this problem more generally, for general hypothesis classes $\mathcal{H}$ and general values $b\geq 0$. We define the notion of the robust hollow star number of $\mathcal{H}$ (which generalizes the standard hollow star number), and show that it precisely characterizes the worst-case size of the smallest certificate achievable, and analyze its size for natural classes. We also consider worst-case distributional bounds on certificate size, as well as distribution-dependent bounds that we show tightly control the sample size needed to get a certificate for any given test example. In particular, we define a notion of the certificate coefficient $\varepsilon_{x}$ of an example $x$ with respect to a data distribution $\mathcal{D}$ and target function $h^{\star}$, and prove matching upper and lower bounds on sample size as a function of $\varepsilon_{x}$, $b$, and the VC dimension $d$ of $\mathcal{H}$.
@@ 1 Introduction
There has been substantial recent interest in explainable AI, (AAES+, 23; DDN+, 23; DVK, 17; RSG, 16; Mil, 19). For example, in a medical setting, if a classifier $h\in\mathcal{H}$ trained on some large dataset $S$ predicts that patient $x$ should get treatment $y$, the patient’s doctor may want an explanation of why. Much of the work in explainable machine learning has focused on decision-tree models, or identifying the most salient features for the prediction made (CS, 95; BS, 96; ZH, 16). In this work, we consider an alternative approach: outputting a subset $S^{\prime}$ of the training set $S$ such that all classifiers $h^{\prime}\in\mathcal{H}$ that agree with $S^{\prime}$ (or that make at most $b$ mistakes on $S^{\prime}$) predict $h^{\prime}(x)=y$, if such an $S^{\prime}$ exists. Such a set $S^{\prime}$ would serve as a proof that $x$ indeed has label $y$ under the assumption that (1) the true target function $h^{\star}$ belongs to $\mathcal{H}$, and (2) the set $S$ contains at most $b$ noisy or corrupted points. For example, if $b=0$ and $\mathcal{H}$ is the family of linear classifiers in $\mathbb{R}^{d}$, and if $x$ lies inside the convex hull of the positive data points in $S$ (and hence every consistent linear classifier labels $x$ as positive), then Carathéodory’s theorem states that $x$ in fact lies inside the convex hull of $d+1$ of those points. So, a set $S^{\prime}$ of size $d+1$ could be released as an explanation for a positive prediction, and a proof of its correctness under the assumption of realizability.
We aim to consider such explanations for general families $\mathcal{H}$ and general values of $b$. Our work is inspired by BBHS (22) who propose the notion of robustly reliable classifiers that, given an example $x$, output both a label $y$ and a value $b$ with the guarantee that any $h^{\prime}\in\mathcal{H}$ with $h^{\prime}(x)\neq y$ makes strictly more than $b$ mistakes on $S$ (where $b<0$ if $x$ is not in the agreement region of the version space). Our work can be viewed as investigating the shortest proof that can be provided for such a guarantee.
@@ 1.1 Main Contributions
    We formalize the notion of a robust certificate: a subset $S^{\prime}$ of the training data that serves as a proof that a given example $x$ must have label $y$ if the target function belongs to a given class $\mathcal{H}$ and the training set has at most $b$ noisy or corrupted points. To analyze this, we define the notion of the robust hollow star number of $\mathcal{H}$, which generalizes the standard hollow star number (BHMZ, 20), and show that it precisely characterizes the worst-case size of the smallest certificate achievable for a class $\mathcal{H}$, and analyze its size for natural classes.
@@
    We examine worst-case distributional bounds on certificate size, showing that in this case, one can achieve tight bounds on certificates achievable from a finite sample in terms of the (standard) hollow star number of BHMZ (20).
@@ 1.2 Context and Related Work
Explainable ML research largely focuses on decision trees or key predictive features. In these approaches, a certificate for an instance $x$ corresponds to the root-to-leaf path of $x$ in the tree. A widely studied method for explaining a black-box model $f$ involves first learning a decision tree $T$ that closely approximates $f$. Once this surrogate decision tree $T$ is obtained, a certificate for any instance $x$ can be generated by retrieving its associated path in $T$ (CS, 95; BS, 96; ZH, 16). BLT (21) proposed algorithms for implicitly learning the surrogate decision trees that approximate the target function, with provable performance guarantees under the uniform distribution. RSG (18) were the first to introduce certificates (which they term anchors) providing high precision by identifying a minimal set of rules that anchor a prediction, ensuring that the output remains stable under small perturbations. They provide an efficient heuristic, based on greedy search, for finding such high-precision certificates. BKLT (22) investigates the minimum number of queries required to certify a monotone function’s prediction at a given point. They define as certificate a subset of input coordinates such that fixing these guarantees the function’s value remains unchanged; GM further investigate this. There is also an increasing interest in hybrid models that are partially interpretable as studied by FLA (23); FLMM (24).
Another related line of work considers learning guarantees in the face of malicious noise (KL, 88; BH, 21), such as the breakthrough work of ABL (17) on guarantees for learning linear separators under malicious noise, as well as work on instance-targeted data-poisoning attacks (BNS+, 06). SMK+ (18) and SHN+ (18) demonstrated empirically that data-poisoning attacks can be highly effective, even when the adversary only adds correctly-labeled data to the training set. These targeted attacks have attracted attention in recent years due to their potential to compromise the trustworthiness of systems (GFH+, 21; MKSKRJ, 15; CLL+, 17).
A key concern here is: when can predictions be trusted under such attacks? Most theoretical work on this question has focused on certifying stability of predictions under small changes to the training set (GKM, 21; LF, 21). However, recently, BBHS (22); BHPS (23); BS (24) consider certifying the actual correctness of predictions, under assumptions of realizability and bounded adversarial power. Our work can be viewed as studying the shortest proof that can be provided for such a guarantee.
@@ 2.1 Notation
The input space is denoted $\mathcal{X}$, and the label space $\{+1,-1\}$. A hypothesis class $\mathcal{H}$ is a subset of $\{-1,+1\}^{\mathcal{X}}$. A sequence $S=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ is realizable by $\mathcal{H}$ if $\exists h\in\mathcal{H}$ such that $\sum_{i=1}^{n}\mathds{1}[h(x_{i})\neq y_{i}]=0$. We use $[n]$ to denote the first $n$ natural numbers, $n]=\{1,2,\dots,n\}$.
@@ 2.2 Formal Setting
When $b=0$, we refer to the 0-robust agreement region simply as the agreement region.
@@
Our setting assumes that we are given a $b$-robustly realizable training dataset $S$, together with a test point $x$ which satisfies that $(x,y)$ belongs to the $b$-robust agreement region of $S$, for some $y\in\{+1,-1\}$. Namely, $S$ is itself a $b$-robust certificate for (x,y)(x,y). [^1] The problem of obtaining small-size subsets of the training data that can serve as good-faith certificates for test-time predictions only makes sense if the test point is in the agreement region of the training data to start with. Our primary objective is to analyze and obtain the smallest $S^{\prime}\subseteq S$ such that $S^{\prime}$ continues to be a $b$-robust certificate for $(x,y)$.
@@ 3 Worst Case Bounds on Certificate Size
Consider the class of $d$-dimensional halfspaces passing through the origin, i.e., $\mathcal{X}=\mathbb{R}^{d}$, $\mathcal{H}=\{x\mapsto\operatorname{sign}(w^{T}x):w\in\mathbb{R}^{d}\}$. Suppose we are given a training dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{H}$. Let $S_{+}=\{x_{i}:i\in[n],y_{i}=+1\}$, $S_{-}=\{-x_{i}:i\in[n],y_{i}=-1\}$, and let $S_{+}\cup S_{-}=\{z_{1},\ldots,z_{n}\}$ for convenience. Note that we negate a negatively labeled $x_{i}$ in $S$ to form $S_{-}$. A test point $(x,+1)$ belongs to the agreement region of $S$ if and only if $x\in\operatorname{Cone}(S_{+}\cup S_{-})$, where $\operatorname{Cone}(\cdot)$ denotes the conic hull. To see this, suppose first that $x\in\operatorname{Cone}(S_{+}\cup S_{-})$. This means that $x=\sum_{i=1}^{n}\alpha_{i}z_{i}$, where $\alpha_{i}\geq 0,\forall i$. Let $w\in\mathbb{R}^{d}$ represent any halfspace that labels all the examples in $S$ correctly. Then, observe that for every $z_{i}\in S_{+}\cup S_{-}$, $w^{T}z_{i}\geq 0$. But this also means that $w^{T}x=\sum_{i=1}^{n}\alpha_{i}\cdot w^{T}z_{i}\geq 0$. Thus, $(x,+1)$ belongs to the agreement region of $S$. In the other direction, suppose that $x\notin\operatorname{Cone}(S_{+}\cup S_{-})$. Then, because $\operatorname{Cone}(S_{+}\cup S_{-})$ is a closed convex set, the separating hyperplane theorem (e.g., see Theorem 1 in Nac (18)) implies the existence of $w\in\mathbb{R}^{d}$, such that $w^{T}x<0$, and $w^{T}y\geq 0$ for $y\in\operatorname{Cone}(S_{+}\cup S_{-})$. In particular, this means that $w$ labels all examples in $S$ correctly. However, $w^{T}x<0$, so $(x,+1)$ is not in the agreement region of $S$.
So, consider a test point $(x,+1)$ that belongs to the agreement region of $S$. By the preceding argument, $x\in\operatorname{Cone}(S_{+}\cup S_{-})$. But then, by Carathéodory’s Theorem, $x$ can be written as a conic combination of at most $d$ points [^2] There are hardness results on computing the “Carathéodory number” of a point set (BH, 20). from $S_{+}\cup S_{-}$, which implies that $(x,+1)$ is in the agreement region of these points. Thus, we can use this subset of $S_{+}\cup S_{-}$ (together with labels), which has size at most $d$, as a certificate for $(x,+1)$.
@@
Consider the class of singletons on a domain of size $n$, i.e., $\mathcal{X}=[n]$, $\mathcal{H}=\{x\mapsto(-1)^{1-\mathds{1}[x=i]}:i\in[n]\}$. The VC dimension of this class is 1. Suppose that we are given a training dataset $S=\{(1,-1),(2,-1),\ldots,(n-1,-1)\}$. Observe that the test point $(n,+1)$ is in the agreement region of $S$, because the only hypothesis in $\mathcal{H}$ that labels $S$ in the given way labels $n$ positively. Note however, that the test point is not in the agreement region of any proper subset of $S$. We must therefore necessarily provide all the n1n-1 points in $S$ to certify a positive label on $n$.
The instance in the above example has a curious property: as long as there remains a single point in the domain whose label we haven’t observed, it is not possible to completely determine the label of the test point. This, however, is the defining property of another combinatorial quantity in learning theory, known as the hollow star number (BHMZ, 20). For example, the hollow star number is known to lower bound the sample complexity of proper PAC learners (BHMZ, 20, Theorem 10). Indeed, as we show ahead, the hollow star number ends up also being the relevant quantity that characterizes worst-case minimum certificate sizes. First, we define a slightly more generalized version of the hollow star number, which allows us to handle $b\geq 0$ corruptions.
@@
A sequence $T=\{(x_{1},y_{1}),\dots,(x_{s_{b}},y_{s_{b}})\}$ is a $b$-robust hollow star for $\mathcal{H}$, if there exist integer weights $w_{1},\dots,w_{s_{b}}$ and $i^{*}\in[s_{b}]$ such that:
@@
    For any $i\in[s_{b}]$, $T\setminus\{(x_{i},y_{i},w_{i})\}$ is $b$-robustly realizable by $\mathcal{H}$.
The size $s_{b}$ of the largest $b$-robust hollow star is the $b$-robust hollow star number of $\mathcal{H}$.
@@
We observe that setting $b=0$ in the above definition recovers the standard definition (e.g., Definition 3 in BHMZ (20)) of the hollow star number. We also note that repeats are allowed in the sequence $T$ (i.e., there might be $i\neq j$ where $x_{i}=x_{j}$).
@@
For $b\geq 0$, let $s_{b}$ be the $b$-robust hollow star number of $\mathcal{H}$. Then, $s_{b}\geq(b+1)(s_{0}-1)+1$.
@@
Let $\mathcal{H}$ be a hypothesis class that has $b$-robust hollow star number $s_{b}$, $S$ be a training dataset that is $b$-robustly realizable by $\mathcal{H}$, and $x$ be a test point such that for some $y\in\{-1,1\}$, $(x,y)$ is in the $b$-robust agreement region of $S$. Then, there exists a $b$-robust certificate $S^{\prime}\subseteq S$ for $(x,y)$ that has size at most $s_{b}-1$. Furthermore, there exists a training dataset $S$ of size $s_{b}-1$ that is $b$-robustly realizable by $\mathcal{H}$, test point $x$ and test label $y\in\{-1,1\}$ which satisfy that $(x,y)$ belongs to the $b$-robust agreement region of $S$, such that no proper subsequence of $S$ is a $b$-robust certificate for $(x,y)$.
@@
Consider $S^{\prime}\subseteq S$ to be the smallest subset of $S$ that is a $b$-robust certificate for (x,y)(x,y)—let $S^{\prime}=\{(x_{1},y_{1}),\dots,(x_{k},y_{k})\}$. Note that $S^{\prime}$ is a minimal $b$-robust certificate for $(x,y)$. We argue that $\{(x_{1},y_{1}),\dots,(x_{k},y_{k}),(x,1-y)\}$ is a $b$-robust hollow star, which implies $k+1\leq s_{b}$.
Consider the weighted sequence $T=\{(x_{1},y_{1},1),\dots,(x_{k},y_{k},1),(x,1-y,b+1)\}$. Now consider any $h\in\mathcal{H}$. If $h(x)=y$, then the weighted error of $h$ on $x$ is $b+1$, and $h$ does not $b$-robustly realize $T$. Otherwise, we have that h(x)=1yh(x)=1-y. In this case, it must be the case that $\sum_{i=1}^{k}\mathds{1}[h(x_{i})\neq y_{i}]>b$, otherwise $S^{\prime}$ would not be a $b$-robust certificate for $(x,y)$. Summarily, the weighted sequence $T$ is not $b$-robustly realizable by any $h\in\mathcal{H}$.
Now, consider $T^{\prime}=T\setminus\{(x,1-y,b+1)\}$. Since $S$ is $b$-robustly realizable by $\mathcal{H}$, so is $S^{\prime}$, which implies that the weighted sequence $T^{\prime}$ is also $b$-robustly realizable by $\mathcal{H}$. Now, consider $T^{\prime}=T\setminus\{(x_{i},y_{i},1)\}$ for any $i\in[k]$. Then, since $S^{\prime}$ is a minimal $b$-robust certificate for $(x,y)$, there must exist some $h\in\mathcal{H}$ such that $\sum_{j=1,j\neq i}^{k}\mathds{1}[h(x_{j})\neq y_{j}]\leq b$ and h(x)=1yh(x)=1-y. Otherwise, $S^{\prime}\setminus\{(x_{i},y_{i})\}$ would be a smaller $b$-robust certificate for $(x,y)$. Consequently, such an $h$ $b$-robustly realizes the weighted sequence $T^{\prime}$, and we are done.